One game at the amusement park has you push a puck up a long, frictionless ramp. You win a stuffed animal if the puck, at its highest point, comes to within 10 cm of the end of the ramp without going off. You give the puck a push, releasing it with a speed of 5.0 m/s when it is 9.3 m from the end of the ramp. The pucks speed after traveling 3.0 m is 4.0 Part A How far is it from the end when it stops?

Part A: How far is it from the end when it stops?
Answer #1

use kinematics to find the acceleration

v2 - u2 = 2ad

42 - 52 = 2*a*3

a = -1.5 m/s2

Now,

distance covered before puck finally stops

02 - 52 = 2*(-1.5)d

d = 8.333 m

so,

distance left = 9.3 - 8.333

distance = 0.9666 m